40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 15233     Entry time: Thu Feb 27 22:45:40 2020
Author: gautam 
Type: Update 
Category: ALS 
Subject: ALS noise high 

There was some UNELOGGED work at EX today. The DFD outputs were also hijacked for loss measurement. Unclear who the culprit was, but there is now a broad noise bump centered around ~180 Hz in the ALS X noise curve, which certainly wasn't there yesterday. Maybe let's keep the few working systems working, it is annoying to have to deal with these auxiliary issues every night. I'll push ahead with locking, hopefully the ALS noise is "good enough".

Attachment 1: ALSnoise.pdf  75 kB  | Hide | Hide all
ALSnoise.pdf
ELOG V3.1.3-